What is the free cash flow of a firm with revenues of $343 million, operating profit margin of 36%, tax rate of 31%, depreciation and amortization expense of $23 million, capital expenditures of $37 million, acquisition expenses of $6 million and change in net working capital of $17 million? Answer in millions, rounded to one decimal place (e.g., $245.63 = 245.6).
Answer #1

Operating profit = 343 * 36% = 123.48

Operating cash flow = EBIT + Depreciation - Taxes

= 123.48 + 23 - ( 123.48 *31%)

= 108.2012

Free cash flow = Operating cash flow - Net capital spending - Change in net working capital

= 108. 2012 - 37 -6 - 17

= 48.2012

Ans: $48.2 [ rounded off]